Grade 11Algebra

Show that :
1 + (23/2! ) + (33/3! ) + (43/4! ) + …. = 5e

Nandish Panchal
7 Years agoGrade 11

1 Answer

Aditya Gupta
7 Years ago
nth term is n^3/n!
= n^2/(n-1)!
= 1/(n-1)!+(n+1)/(n-2)!
= 1/(n-1)!+1/(n-3)!+ 3/(n-2)!
now summation upto infinity is
= e+e+3e
=5e